OpenStudy (anonymous):

Evaluate (nm)3, when n = 5 and m = 3

OpenStudy (unklerhaukus):

is that \((nm)^3\)?

OpenStudy (unklerhaukus):

\((ab)^c=a^c\times b^c\)